A fullback preparing to carry the football starts from rest and accelerates straight ahead. He is handed the ball just before he reaches the line of scrimmage. Assume that the fullback accelerates uniformly (even during the handoff), reaching the line with a velocity of 7.56 m/s. If he takes 1.04 s to reach the line, how far behind it did he start?
Answer #1

here,

final speed of ball , v = 7.56 m/s

time taken , t = 1.04 s

let the accelration be a

using first equation of motion

v = u + a * t

7.56 = 0 + a * 1.04

solving for a

a = 7.27 m/s^2

the distance travelled while accelrating , s = u * t + 0.5 * a * t^2

s = 0 + 0.5 * 7.27 * 1.04^2 m

s = 3.93 m